OpenStudy (anonymous):

Express in terms of i: (square root of -12) - (square root of -3)

OpenStudy (anonymous):

= sqrt12 *i - sqrt3*i = 2sqrt3 *i - sqrt3 * i = sqrt3*i

OpenStudy (anonymous):

because sqrt12 = sqrt(4*3) = sqrt4 * sqrt3 = 2sqrt3